OpenStudy (anonymous):

Find the slope of the ling that passes through each pair of points. (-3,-2) & (4,-2).

OpenStudy (anonymous):

To find slope, you would use the formula... \[y_2-y_1 \over x_2-x_1\]

OpenStudy (anonymous):

0/1 which would be a slope of 0

OpenStudy (anonymous):

Err I mean 0/7

OpenStudy (anonymous):

would still be a slope of 0

OpenStudy (anonymous):

So using that formula you get:\[\frac{-2 - -2}{-3-4} = \frac{0}{-7} = 0\]

OpenStudy (anonymous):

For future reference, if you wound up with a number over 0 say \[7 \over 0\] then slope would be undefined

OpenStudy (anonymous):

Oh okay! thank you.

OpenStudy (anonymous):

You're very welcome :)

OpenStudy (vijay):

thanks for delete my answer victoria, gud bye

OpenStudy (anonymous):

What? I cant delete your answer, only you can... But alrighty, buhbye.